How many electrons are transferred in the redox reaction taking place in basic solution:

Cr(s) + ClO4-(aq) --> ClO3-(aq) + Cr3+(aq)?
Answer #1

let us calculate the oxidation state of Cl in ClO4-

let oxidation state be x

charge over each oxide is -2

so we can right

x+(4*-2)=-1 as the overall charge is -1

x-8=-1

x=-1+8=7

so oxidation state of Cl is +7

let us calculate the oxidation state of lr in CrO3-1

let oxidation state be y

charge over each oxide is -2

so we can right

y+(3*-2)=-1 as the molecule is negatively charge or overall charge is -1

y-6=-1

y=-1+6=5

so oxidation state of Cl is +5

we see oxidation state of Cl goes from +7 to +5 there fore it gain of electron and is getting reduced

the change oxidation state for chromium # from0 so it is losing electron and is getting oxidised

half cell oxidation

Cr\rightarrow Cr3+

charge on left side is 0 and right side is +3 so if we add +3e on right side we can balance the charge

we get

Cr\rightarrow Cr3+ +3e equation 1

half cell reduction

ClO4-\rightarrowClO3-

we see the number of oxygen atom on left side is 4 and right side is 3

so in basic condition

for every excess of oxygen 1 H2O added to side which have more oxygen atom and 2OH- molecule on the side which is deficient in it

therefore in this case 1H2O added to left side  and 2OH- molecule on the right side to balance O atoms

so we get

ClO4- +1H2O \rightarrowClO3- +2OH-

we see all atoms are balance (mass balanced equation)

charge on left side is -1 and right side is -2 +(-1)=-3  so if we add +2e on left side we can balance the charge

we get

ClO4- +1H2O + 2e \rightarrowClO3- +2OH- equation 2

Cr\rightarrow Cr3+ +3e equation 1

now multiply equation 1 with 2 and equation 2 with 3 to make the electron count same

3ClO4- +3H2O + 6e \rightarrow3ClO3- +6OH- equation 4

2Cr\rightarrow 2Cr3+ +6e equation 3

net equation adding equation 3 and 4

3ClO4- +3H2O +2Cr \rightarrow3ClO3- +6OH +2Cr3+

thus 6 electrons are transferred in the redox reaction taking place in basic solution